# Interactive Rendering of Relightable and Animatable Gaussian Avatars

### v1

@@ 1 Introduction
To address the aforementioned issues, we propose a new Gaussian representation for relighting avatars during animation. We first define a body mesh and Gaussians in the canonical space, as shown in Figure 1, and the mesh vertices and Gaussians can be animated to the posed space via a body model (i.e., SMPL [26]). The Gaussian properties are interpolated from the mesh vertex properties, which include basic Gaussian properties and material properties. During Gaussian optimization, we optimize the basic Gaussian properties, the material properties and the environment light simultaneously so that the new shading color could be computed under novel illuminations.
@@ 2.1 Human Avatar
2.2 Inverse Rendering
Inverse rendering aims to recover the material of an object from multi-view photos, enabling it to be relighted under novel lighting conditions. Many studies use NeRF to solve the inverse rendering problem [44, 45, 46, 47, 48, 49, 50]. They typically encode geometric and material information into the radiance field, learn the outgoing radiance from captured images, and optimize the materials to decouple the environmental light and materials. [46, 47, 49] use SDF to reconstruct and separate explicit geometric information for better material decoupling.
@@ 2.3 Human Relighting
Relighting4D [14] proposes a method to reconstruct human geometry, recover human materials under unknown environmental lighting, and perform relighting. This method assigns latent features to each frame, which are then used in the MLP for modeling appearance and occlusion maps. Therefore, Relighting4D cannot transfer to novel poses. Sun et al.[65] uses inverse mapping to transform points from the observation space to the canonical space to obtain the material attributes, which are then used for shading color calculation. RANA [66] fits the person via the SMPL+D model and trains networks to further refine albedo and normal. However, neither Sun et al. or RANA models the visibility, therefore the relighted body lacks shadows. To model shadows, many methods provide their unique solutions. RelightableAvatar-Lin [15] trains the part-wise MLP to estimate visibility under novel poses. Bolanos et al. [10] uses a Gaussian density model as an approximation to the NeRF’s density field for easier visibility calculation. RelightableAvatar-Xu [13] proposes Hierarchical Distance Query on the SDF field for sphere tracing, and further utilizes Distance Field Soft Shadow (DFSS) for soft visibility. IntrinsicAvatar [11] implements the body inverse rendering by explicit ray tracing and introducing the secondary shading effects, thus modeling accurate shadows naturally. The above methods successfully model the shadows generated by body occlusion, achieving realistic relighting results. However, these methods are based on neural rendering, which requires multiple samplings and therefore slowing down the speed. They all require several seconds for an image, restricting their applications in non performance critical applications. By using 3DGS, Li et al. [12] achieves excellent relighting results with unique body details under novel poses. However, it is trained under relatively dense viewpoints and relies on ray tracing to calculate visibility, which still requires several seconds to render a picture. We noticed that just two days before the paper submission, a concurrent work related to relightable Gaussian avatars [67] was published online. This work uses a UNet to model the pose-dependent material fields, and Monte-Carlo sampling to compute the outgoing radiance. It is trained using videos from 16 viewpoints, and since the code has not been released yet, it is still unclear whether it can achieve training with sparse-view or monocular videos as ours.
@@ 3.1 Relightable Gaussian Avatar
* represents one of the attributes $\{\mathbf{r},\mathbf{s},\mathbf{n},\mathbf{a},\gamma,p,\mathbf{w},\Delta\mathbf{x},\mathbf{v}\}$. $*^{j}$ and $*_{g}^{i}$ denote the attributes of the $j$th vertex and the $i$th Gaussian, respectively. $\mathcal{S}^{i}_{g}$ contains the indices of the $K_{g}$ nearest vertices to the Gaussian $\mathbf{x}_{g}^{i}$. $u_{j}(\mathbf{x}^{i})=1/\|\mathbf{x}_{g}^{i}-\mathbf{x}^{j}\|_{2}$ is the interpolation weight. We find it’s sufficient to model the space-varying high-frequency attributes by directly interpolating vertices if the vertex density is high. We always set the opacity of the Gaussians to $1$, so there is no need to model the opacity attribute.
@@
For the environment map, we use the same design as [49, 14, 13], which define the map as light probes $L\in\mathbb{R}^{16\times 32\times 3}$ with 512 discrete area lights. Figure 1 displays the spherical format of the environment map. At the posed space, to calculate the shading of a sample Gaussian at a given position $\mathbf{\bar{x}}_{gd}$ viewed from a certain direction $w_{o}$, we integrate the rendering equation by explicitly summing up the discrete light probes,
@@ 3.3 Training
Method Albedo Relighting (Training poses) Relighting (Novel poses)
Ours 34.2106 0.9418 0.1705 36.2018 0.8639 0.0560 28.7703 0.8304 0.0740
RA-Lin [15] 29.6458 0.9287 0.1775 32.3223 0.8501 0.0661 28.6252 0.8290 0.0775
RA-Xu [13] 31.4815 0.9346 0.1807 33.1538 0.8526 0.0694 28.4215 0.8258 0.0828
IA [11] 27.0414 0.9165 0.1993 30.0939 0.8365 0.0816 28.3806 0.8256 0.0880
R4D [14] 23.4056 0.8920 0.2228 27.7852 0.8179 0.1072 24.5498 0.7989 0.1178
@@ 3.5 Implementation Details
After training, the trained avatar can be relighted by replacing the trained environment map $L$ with a novel environment map $L^{\prime}$ and animated with novel poses. The visibility is computed on the fly during testing. In order to make the model generalize better to novel poses, for a vertex $\mathbf{x}^{i}$, we calculate the average displacement $\Delta\mathbf{\tilde{x}}^{i}=\frac{1}{T}\sum_{t=1}^{T}d(\mathbf{x}^{i},\theta^{t})$ across all the training poses. During novel pose testing, we use the average displacement $\Delta\mathbf{\tilde{x}}^{i}$ to replace the vertex’s original displacement attribute $\Delta\mathbf{x}^{i}$.
@@
Figure 2: Qualitative comparison with RA-Lin [15], RA-Xu [13], IA [11] and R4D [14]. We show the albedo and the relighting results under training and novel poses on both synthetic data (jody, rendered at test viewpoints) and real data (ZJU-377 and male-3-casual, rendered at training viewpoints). Compared to the baselines, our method can achieve finer body details (jody’s leggings, ZJU-377’s face and male-3-casual’s jeans) and the specular effects that are closest to the ground truth (jody’s leggings).
@@ 4 Evaluation
In this section, we conduct experiments on multiple datasets (Section 4.1) and introduce the metrics to validate the results (Section 4.2). Compared to baselines, our results outperform others in achieving human body relighting (Section 4.3). We conduct ablation studies that validate the effectiveness of several designs (Section 4.4). We also show the appearance editing ability of our method (Section 4.5). In terms of rendering efficiency, our method surpasses previous works, achieving rendering at interactive frame rates (Section 4.6). Additionally, we show the interactive relighting results in supplementary video.
@@ 4.2 Metrics
For quantitative evaluations, we use Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM), and Learned Perceptual Image Patch Similarity (LPIPS)[73] as metrics. On synthetic data, we render the albedo and the relighting results from both training poses and novel poses for metric computation. Our method can render at high resolution (1K $\times$ 1K). However, for fair comparison, in Section 4.3 and Section 4.6, all the comparison experiments are trained and rendered at a resolution of 500 $\times$ 500. In other experiments, we train and test our method at a resolution of 1K. Our metrics are calculated on the entire image, including the black background. Similar to [11], we compute per-channel scaling factor to align the albedo and rendered images with the ground truth, addressing the ambiguity issue of inverse rendering between different methods.
@@ 4.3 Comparison
Baselines We compare our work with RelightableAvatar-Lin (RA-lin [15]), RelightableAvatar-Xu (RA-Xu [13]), IntrinsicAvatar (IA [11]), and Relighting4D (R4D [14]). To ensure fair comparison, we compare our method where the data and source code are available. Some baselines [14, 11] are originally designed for monocular videos. We adapt them for a multi-view setting. Other baselines such as [12, 10, 65, 66, 67] are not employed due to the lack of data or code.
Results Figure 2 shows the comparison. We present the results under both synthetic and real datasets. The images are all adjusted by per-channel scaling to make the brightness similar for different methods. R4D [14] assigns latent codes to each frame to model the appearance, therefore the reconstructed human body naturally fails to generalize to novel poses. For IA [11], the results appear relatively blurry, because IA doesn’t model pose-dependent non-rigid deformation for the animated human. For RA-Xu [13], its design of the signed distance function tends to create hollow artifacts under the armpits (Figure 2, red box). Overall, although our results seem roughly similar to those of RA-Lin [15] and RA-Xu [13], our results perform better in preserving the details. Compared with RA-Lin and RA-Xu, our method achieves clearer details on the stripes of jody’s leggings, the face of ZJU-377 and the jeans of male-3-casual (Figure 2, blue box). Our method also produces specular results closer to ground truth on jody’s leggings. Table I shows the quantitative results between our method and the baselines. Our method outperforms other methods under various metrics. We note that in the novel pose test, due to some misalignments between the posed human body and the ground truth, there will be a noticeable drop in metrics, as illustrated in [15] as well. The comparison results show that our method can better recover materials and render superior relighting results.
@@ 4.4 Ablation Study
Our densification For the case of male-3-casual, the character maintains an A-pose in the training data and the underarm area is not visible to the camera for all the frames. As a result, Gaussians barely move and the view-space positional gradients are small, thus the original 3DGS’s densification doesn’t increase the number of Gaussians under the arm. We add Gaussians to areas where the Gaussian density is low, such as the armpits, to prevent holes when the avater is animated under a novel pose (See Figure 3).
@@ 4.6 Speed
RA-Lin [15] RA-Xu [13] IA [11] R4D [14] Ours
Training 44h 33h 3.5h 49h 55min (+6h, SDF mesh)
Render 15.0s 4.9s 5.5s 4.0s 0.145s (0.021s, no visibility)
@@
4.7 Limitation
Similar to previous methods [15, 13, 11], we only learn consistent materials from the video frames, so the method cannot model pose-dependent wrinkle deformation. Our method also cannot reconstruct loose clothing, like dress. Our method uses [15] to obtain body mesh, which utilizes MLP to model the SDF and is slow to train. Some methods based on iNGP [9] could acquire the mesh faster.

### v2

@@ 1 Introduction
To address the aforementioned issues, we propose a new Gaussian representation for relighting avatars during animation. We first define a body mesh and Gaussians in the canonical space, as shown in Figure 1, and the mesh vertices and Gaussians can be animated to the posed space via a body model (i.e., SMPL [26]).
The Gaussian properties are interpolated from the mesh vertex properties, which include basic Gaussian properties (position, rotation, scale, and opacity) and material properties. During Gaussian optimization, we optimize the basic Gaussian properties, the material properties and the environment light simultaneously so that the new shading color could be computed under novel illuminations.
@@ 2.1 Human Avatar
2.2 Radiance-field-based Inverse Rendering
Inverse rendering aims to recover the material of an object from multi-view photos, enabling it to be relighted under novel lighting conditions. Many studies use NeRF to solve the inverse rendering problem [44, 45, 46, 47, 48, 49, 50]. They typically encode geometric and material information into the radiance field, learn the outgoing radiance from captured images, and optimize the materials to decouple the environmental light and materials. [46, 47, 49] use SDF to reconstruct and separate explicit geometric information for better material decoupling. Zhang et al. [46] can model the indirect illumination. PhySG [51] represents specular BRDFs and environmental illumination using mixtures of spherical Gaussians. NeRD [52] and Neural-PIL [53] can decouple materials using images captured under different illumination conditions. Lyu et al. [54] generate relighting results with global illumination. NeRO [55] can reconstruct the BRDF of reflective objects with strong reflective appearances. Mai et al. [56] use a microfacet reflectance model to recover high-quality materials, geometry, and illumination, while NeMF [57] uses microflake volume to relight complex objects.
@@ 2.3 Human Relighting
Relighting4D [14] proposes a method to reconstruct human geometry, recover human materials under unknown environmental lighting, and perform relighting. This method assigns latent features to each frame, which are then used in the MLP for modeling appearance and occlusion maps. Therefore, Relighting4D cannot transfer to novel poses. Sun et al.[72] uses inverse mapping to transform points from the observation space to the canonical space to obtain the material attributes, which are then used for shading color calculation. RANA [73] fits the person via the SMPL+D model and trains networks to further refine albedo and normal. However, neither Sun et al. or RANA models the visibility, therefore the relighted body lacks shadows. To model shadows, many methods provide their unique solutions. RelightableAvatar-Lin [15] trains the part-wise MLP to estimate visibility under novel poses. Bolanos et al. [10] uses a Gaussian density model as an approximation to the NeRF’s density field for easier visibility calculation. RelightableAvatar-Xu [13] proposes Hierarchical Distance Query on the SDF field for sphere tracing, and further utilizes Distance Field Soft Shadow (DFSS) for soft visibility. IntrinsicAvatar [11] implements the body inverse rendering by explicit ray tracing and introducing the secondary shading effects, thus modeling accurate shadows naturally. The above methods successfully model the shadows generated by body occlusion, achieving realistic relighting results. However, these methods are based on neural fields, which require expensive ray marching with many sample point queries with MLPs, therefore slowing down the speed. They all require several seconds for an image, restricting their applications in non performance critical applications. RGCA [6] is able to create realistic and relightable head avatar, but this method requires to capture the data in a light stage with controllable lights and hundreds of cameras, which is difficult for ordinary users to obtain. By using 3DGS, Li et al. [12] achieves excellent relighting results with unique body details under novel poses. However, it is trained under relatively dense viewpoints and relies on ray tracing to calculate visibility, which still requires several seconds to render a picture. MeshAvatar [74] uses a UNet to model the pose-dependent material fields, and Monte-Carlo sampling to compute the outgoing radiance. However, it cannot recover fine geometry under sparse-view settings, resulting in suboptimal relighting results (see Figure 2).
@@ 3.1 Relightable Gaussian Avatar
* represents one of the attributes $\{\mathbf{r},\mathbf{s},\mathbf{n},\mathbf{a},\gamma,p,\mathbf{w},\Delta\mathbf{x},\mathbf{v}\}$. $*^{j}$ and $*_{g}^{i}$ denote the attributes of the $j$th vertex and the $i$th Gaussian, respectively. $\mathcal{S}^{i}_{g}$ contains the indices of the $K_{g}$ nearest vertices to the Gaussian $\mathbf{x}_{g}^{i}$. $u_{j}(\mathbf{x}^{i})=1/\|\mathbf{x}_{g}^{i}-\mathbf{x}^{j}\|_{2}$ is the interpolation weight. We find it’s sufficient to model the spatially varying high-frequency attributes by directly interpolating vertices if the vertex density is high. We always set the opacity of the Gaussians to $1$, so there is no need to model the opacity attribute.
@@
For the environment map, we use the same design as [49, 14, 13], which define the map as light probes $L\in\mathbb{R}^{16\times 32\times 3}$ with 512 discrete area lights. Figure 1 displays the spherical format of the environment map. In the posed space, to calculate the shading of a sample Gaussian at a given position $\mathbf{\bar{x}}_{gd}$ viewed from a certain direction $w_{o}$, we integrate the rendering equation by explicitly summing up the discrete light probes,
@@ 3.2 Visibility Computation
Discussion with shadow mapping. While our method shares a similar idea with shadow mapping [78] in rasterizing meshes from the light source for fast shadow calculation, there is still a difference. Shadow mapping rasterizes a depth map from the light source and conducts depth testing from viewpoints to check the visibility of a point on the mesh, while our method rasterizes from the light source to obtain a triangle’s visibility and directly uses triangle’s visibility to approximate a point’s visibility.
Method Normal Albedo Relighting (Training poses) Relighting (Novel poses)
Error $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Ours 9.322 34.3489 0.9428 0.1695 36.1635 0.8640 0.0555 28.8396 0.8308 0.0737
RA-Lin [15] 10.155 30.6854 0.9342 0.1766 33.1736 0.8548 0.0634 28.4573 0.8281 0.0774
MeshAvatar [74] 11.685 30.5407 0.9283 0.1859 33.4034 0.8522 0.0724 27.9275 0.8217 0.0878
RA-Xu [13] 10.132 32.0395 0.9365 0.1790 33.5669 0.8544 0.0675 28.3756 0.8258 0.0817
IA [11] 11.774 27.9200 0.9252 0.1839 31.5716 0.8463 0.0693 28.1717 0.8259 0.0807
R4D [14] 24.502 24.9638 0.9026 0.2050 28.9139 0.8306 0.0879 25.4990 0.8081 0.1024
@@ 3.4 Appearance Editing
Figure 2: Qualitative comparison with RA-Lin [15], MeshAvatar [74], RA-Xu [13], IA [11] and R4D [14]. We show the albedo and the relighting results under training and novel poses on both synthetic data (jody, rendered at test viewpoints) and real data (ZJU-377 and male-3-casual, rendered at training viewpoints). Compared to the baselines, our method can achieve finer body details (jody’s leggings, ZJU-377’s face and male-3-casual’s jeans) and the specular effects that are closest to the ground truth (jody’s leggings).
@@ 3.5 Implementation Details
After training, the trained avatar can be relit by replacing the trained environment map $L$ with a novel environment map $L^{\prime}$ and animated with novel poses. The visibility is computed on the fly during testing. In order to make the model generalize better to novel poses, for a vertex $\mathbf{x}^{i}$, we calculate the average displacement $\Delta\mathbf{\tilde{x}}^{i}=\frac{1}{T}\sum_{t=1}^{T}d(\mathbf{x}^{i},\theta^{t})$ across all the training poses. During novel pose testing, we use the average displacement $\Delta\mathbf{\tilde{x}}^{i}$ to replace the vertex’s original displacement attribute $\Delta\mathbf{x}^{i}$.
@@ 4 Evaluation
In this section, we conduct experiments on multiple datasets (Section 4.1) and introduce the metrics to validate the results (Section 4.2). Our approach outperforms baselines in achieving human body relighting (Section 4.3). We conduct ablation studies that validate the effectiveness of several designs (Section 4.4). We also show the appearance editing ability of our method (Section 4.5). In terms of rendering efficiency, our method surpasses previous works, achieving rendering at interactive frame rates (Section 4.6). Additionally, we show the interactive relighting results in supplementary video.
@@ 4.2 Metrics
For quantitative evaluations, we use Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM), and Learned Perceptual Image Patch Similarity (LPIPS)[81] as metrics. We also follow IA [11] and RA-Xu [13] to compute the normal difference (in degrees) between the results and the ground truth. On synthetic data, we render the albedo and the relighting results from both training poses and novel poses for metric computation. Our method can render at high resolution (1K $\times$ 1K). However, for fair comparison, in Section 4.3 and Section 4.6, all the comparison experiments are trained and rendered at a resolution of 500 $\times$ 500. In other experiments, we train and test our method at a resolution of 1K. Our metrics are calculated on the entire image, including the black background. Similar to [11], we compute per-channel scaling factor to align the albedo and rendered images with the ground truth, addressing the ambiguity issue of inverse rendering between different methods.
@@ 4.3 Comparison
Baselines We compare our work with RelightableAvatar-Lin (RA-lin [15]), RelightableAvatar-Xu (RA-Xu [13]), IntrinsicAvatar (IA [11]), Relighting4D (R4D [14]) and MeshAvatar [74]. To ensure fair comparison, we compare our method where the data and source code are available. Some baselines [14, 11] are originally designed for monocular videos. We adapt them for a multi-view setting. Other baselines such as [12, 10, 72, 73, 74] are not employed due to the lack of data or code. To exclude the effect of pose correction for a fair comparison, we use the optimized poses from our method, and disable the pose correction stage for all methods during training.
Results Figure 2 shows the comparison. We present the results under both synthetic and real datasets. The images are all adjusted by per-channel scaling to make the brightness similar for different methods. R4D [14] assigns latent codes to each frame to model the appearance, therefore the reconstructed human body naturally fails to generalize to novel poses. For IA [11], the results appear relatively blurry, because IA doesn’t model pose-dependent non-rigid deformation for the animated human. MeshAvatar [74] reconstructs a non-smooth body surface under the sparse-view setting, resulting in noisy relighting outcomes. For RA-Xu [13], its design of the signed distance function tends to create hollow artifacts under the armpits (Figure 2, red box). Overall, although our results seem roughly similar to those of RA-Lin [15] and RA-Xu [13], our results perform better in preserving the details. Compared with RA-Lin and RA-Xu, our method achieves clearer details on the stripes of jody’s leggings, the face of ZJU-377 and the jeans of male-3-casual (Figure 2, blue box). Our method also produces specular results closer to ground truth on jody’s leggings. Table I shows the quantitative results between our method and the baselines. Our method outperforms other methods under various metrics. We note that in the novel pose test, due to some misalignments between the posed human body and the ground truth, there will be a noticeable drop in metrics, as illustrated in [15] as well. The comparison results show that our method can better recover materials and render superior relighting results.
@@ 4.4 Ablation Study
Our densification For the case of male-3-casual, the character maintains an A-pose in the training data and the underarm area is not visible to the camera for all the frames. As a result, Gaussians barely move and the view-space positional gradients are small, thus the original 3DGS’s densification doesn’t increase the number of Gaussians under the arm. We add Gaussians to areas where the Gaussian density is low, such as the armpits, to prevent holes when the avatar is animated under a novel pose (See Figure 3).
@@ 4.6 Speed
RA-Lin [15] MeshAvatar [74] RA-Xu [13] IA [11] R4D [14] Ours
Training 44h 8.6h 33h 3.5h 49h 55min (+6h, SDF mesh)
Render 15.0s 0.180s 4.9s 5.5s 4.0s 0.145s (0.021s, no visibility)
@@
4.7 Limitation and Future Work
Similar to previous methods [15, 13, 11], our method cannot model pose-dependent wrinkles because the material properties (albedo, roughness, specular tint) do not change with the poses, and we plan to model pose-dependent materials in the future. Our method also cannot reconstruct loose clothing, like dress. Using physics-based simulation to model the geometry of the clothing and optimizing the Gaussians based on the simulated geometry may resolve this issue, which we will leave for future work. Our method uses [15] to obtain body mesh, which utilizes MLP to model the SDF and is slow to train. Some methods based on iNGP [9] could acquire the mesh faster.
@@ 5 Conclusion
Acknowledgment
The authors would like to thank reviewers for their insightful comments. This work is supported by the National Key Research and Development Program of China (No.2022YFF0902302), NSF China (No. 62322209 and No. 62421003), the gift from Adobe Research, the XPLORER PRIZE, and the 100 Talents Program of Zhejiang University.
@@ Appendix B Gaussian Densification Algorithm
Figure 9: Qualitative comparison with RA-Lin [15] and MeshAvatar [74] on ActorsHQ and AvatarRex dataset.
Figure 10: Additional relighting results of our method on ActorsHQ and AvatarRex dataset.
Appendix C Comparison and Results on Other Datasets
We further use ActorsHQ (dataset of HumanRF [83]) and AvatarRex [84] dataset in our experiments to validate our method on higher resolution datasets. ActorsHQ is a high-quality dataset. We select two sequences (actor05, actor07) from ActorsHQ. We use 7 viewpoints and 150 frames for each sequence, with each image approximately 1K resolution. AvatarRex dataset contains full-body multi-view videos. We select two sequences (avatarrex_zzr, avatarrex_zxc) from AvatarRex. We use 6 viewpoints and 120 frames per sequence in our experiments, with each frame approximately 1K resolution.
As these two datasets do not have groundtruth for relighting, we compare our method qualitatively with MeshAvatar [74] and RA-Lin [15] using the above datasets. Figure 9 shows the visual results. MeshAvatar fails to recover smooth body geometry under sparse viewpoints, resulting in noisy relighting outputs. Our relighting results are comparable to those from RA-Lin [15], and outperform MeshAvatar. We also note that the rendering speed for an image of 1K resolution is very slow for RA-Lin (RA-Lin 134.72s vs. ours 0.14s). This is because RA-Lin uses the part-wise MLPs to estimate the visibility, which takes a lot of time.
We also present additional relighting results of our method on ActorsHQ and AvatarRex datasets in Figure 10. All avatars are driven by novel poses. The results demonstrate that our method can reconstruct human avatars from high-resolution datasets like ActorsHQ and AvatarRex, and produce high-quality relighting results under challenging poses.
TABLE V: Quantitative results of different strategies.
Method Albedo Relighting (Training poses) Relighting (Novel poses)
Predict displacement, scale, rotation 34.4069 0.9581 0.1984 36.1664 0.8827 0.1652 28.3171 0.8543 0.1817
Optimize opacity 34.2526 0.9576 0.1983 36.0631 0.8825 0.1647 28.3726 0.8546 0.1816
Appendix D Additional Ablation Study
Predicting displacements, Gaussian rotation offsets and scale offsets. We only predict the displacements in our method. We further follow 4D Gaussian splatting [85] to predict displacements, Gaussian rotation offsets and scale offsets to validate if this design can improve the results. Table V illustrates the results. The quantitative results are comparable to those without predicting the changes of rotation and scaling. We also present a qualitative result, shown in Figure 11. Both images are similar. So we choose to only predict the displacements to reduce the computational burden.
Optimizing opacity. We always set opacity to 1 in our method. We further set the opacity as a learnable parameter and see if this will result in translucent artifacts. Figure 12 (a) shows the opacity map of two designs, and Figure 12 (b) presents the relighting results. As shown in the figure, learnable opacity does not result in translucent artifacts, and the relighting results of optimizing the opacity are almost the same as those from the proposed method. We also present quantitative results of optimizing the opacity in Table V. The quantitative results are also comparable to those without optimizing the opacity. Since optimizing the opacity does not significantly improve the results, we choose to set the opacity to 1.
Figure 11: Qualitative results of predicting displacements, Gaussian rotation offsets and scale offsets.
Figure 12: Qualitative results of optimizing the opacity.